solve. x=(x-2)squareroot of x
Answer by RAY100(1637)

x = (x-2) sqrtx
.
x/ (x-2) = sqrt x
.
square both sides
.
x^2 / (x-2)^2 =x
.
x/ (x-2)^2 =1
.
x= x^2 -4x +4
.
0 = x^2 -5x +4
.
0 = ( x-4)(x-1)
.
x=4, ( 1,,,,,extraneous answer)
.
checking
4= (4-2) sqrt4 = 2*2 ,,,,ok
1=(1-2) sqrt1 = -1(1) =-1,,,,therefore extraneous answer